OpenStudy (anonymous):

A plastic right cylinder with closed ends is to hold V cubic feet. If there is no waste in construction, find the ratio between the height and diameter that results in the minimum use of materials.

OpenStudy (anonymous):

\[V =\pi r^2 h\] \[S.A. = 2 \pi r^2 + 2 \pi r h\]